The central atom in the most reasonable Lewis structure for BrF3 contains three bonding pairs and

four lone pairs.
two lone pairs.
three lone pairs.
one lone electron.
no lone pairs or lone electrons.
one lone pair.
Answer #1

Answer :-

In BrF3 , the central atom is Br which is 17 group element. It has 7 valence electrons of which 3 electrons used in forming bond with F atoms and remained electrons = 7-3 = 4

So lone pairs = 4/2 = 2

Two lone pairs

It is correct answer.